The sum of the digits of a two-digit number is 14. The first digit is 4 less than twice the second digit. What is the number?
Answer by prateekagrawal(56)

here let the number be xy.
then x+y=14 --------(1)
and
x=2y-4 -------------(2) (first digit is 4 less than twice the second digit)
.
so, from (2) put the value of x in (1):
.
2y-4 + y= 14
3y -4=14
3y=18
y=6
so, x+6=14
so, x= 8
.
so, number is 86.